A 2 sided coin is unfairly weighted so when its tossed, the probability that heads will result is twice the probability that tails will result. If the coin is tossed 3x, what is the probability that tails will result exactly 2 times?

Answer by stanbon(75887)

A 2 sided coin is unfairly weighted so when its tossed, the probability that heads will result is twice the probability that tails will result. If the coin is tossed 3x, what is the probability that tails will result exactly 2 times?
----
P(head) = 2/3
P(tail) = 1/3
----
P(2 head in three tosses) = 3C2(2/3)^2(1/3) = 3*(4/9)(1/3) = 12/27
-------------
Cheers,
Stan H.
-------------